# Topology-induced symmetry breaking: a demonstration in antiferromagnetic magnons on a Möbius strip

### v3

@@ Abstract
We propose a mechanism of topology-induced symmetry breaking, where certain local symmetry preserved by the Hamiltonian is broken in the excited eigenstates due to the nontrivial boundary condition. As a demonstration, we study magnon excitations on a Möbius strip comprising of two antiferromagnetically coupled spin chains. Even under a simple Hamiltonian respecting local rotational symmetry and without considering curvature effects, magnons exhibit linear polarization of the Néel vector devoid of chirality and form two non-degenerate branches that cannot be smoothly connected to or be decomposed by the circularly-polarized magnons typically seen in antiferromagnets. One branch supports standing-wave formation on the Möbius strip while the other does not, owing to its spectral shift incurred by the boundary condition. Our findings showcase the significant influence of real-space topology on the physical nature of quasiparticles.
@@
Introduction. In solid-state systems, physical properties of quasiparticles are believed to be direct manifestations of symmetry and interactions, whereas the subtle impact of real-space topology remains elusive. Prevailing studies often adopt periodic boundary conditions (PBCs) on the quasiparticles [1], for which the system becomes topologically equivalent to a circle, a torus, or a 3D-torus depending on the dimensionality. However, there exist exotic structures such as Möbius strips and Klein bottles that do not conform with the PBCs. Concerning the physical behavior of quasiparticles on such an object, it is tempting to ask: what are the physical implications of the non-trivial boundary conditions?
@@
By topological nature, a Möbius strip is non-orientable with a single surface and a single edge, hence precluding the application of any ordinary PBC. Recently, elementary excitations residing on Möbius strips have aroused increasing theoretical attentions [2, 3, 4, 5, 6, 7, 8, 9]. Experimentally, Möbius strips have been realized in a wide range of systems such as molecules [10, 11], single crystals [12], resonators [13, 14, 15], and optical cavities [16, 17], fertilizing a vibrant arena for exploring new physics emerging from the Möbius topology. However, most of these studies focused on the local curvature effects of the strip based on continuous geometry. It is far from clear if there are any residual consequences arising only from the Möbius boundary condition when spatial curvature is discarded.
@@
In this Letter, we propose a hitherto unknown mechanism, dubbed topology-induced symmetry breaking (TISB), which unravels the extraordinary behavior of quasiparticles enabled exclusively by the topological boundary conditions (TBCs) but has nothing to do with any curvature effects. TISB refers to a situation that certain local symmetry respected by the Hamiltonian is explicitly broken in the eigenspace of quasiparticles. While this new mechanism could entail profound consequences in various systems with non-trivial real-space topology, here we demonstrate its manifestation in magnonic excitations on a Möbius strip.
As illustrated in Fig. 1(a), we consider a nano-ribbon composed of two ferromagnetic spin chains oppositely aligned, forming an effective antiferromagnetic (AFM) system involving two sub-lattices. The nano-ribbon can be wrapped into a Möbius strip using two distinct ways of twisting based on Fig. 1(b) and (c). To motivate the following discussions, we first make a critical observation: although the ground-state AFM spin configuration is compatible with the Möbius TBC, the magnons excitations with circular polarizations are inherently irreconcilable with the system topology. Specifically, when only the exchange interactions and the easy-axis anisotropy are considered, the system respects the $O(2)$ rotational symmetry in the spin space. Consequently, the magnon eigenmodes disregarding the TBC are circularly polarized, exhibiting either left-handed or right-handed chirality for both spin species [18, 19, 20]. For example, in the left-handed mode, $\bm{S}_{A}$ and $\bm{S}_{B}$ both rotate clockwisely, but $\bm{S}_{A}$ has a larger oscillation amplitude, which is indicated by the black dashed arrows. As illustrated in Fig. 1(b) and (c), however, imposing the Möbius TBC by connecting the $1$ and $N$ sites with A and B swapped will inevitably disrupt both the chirality and the amplitude of the spin precessions. Therefore, the magnon solutions that are commensurate with the Möbius topology must be fundamentally different from the circularly-polarized magnons widely known for collinear AFM materials.
Model and Results. Let us start with the AFM spin configuration depicted in Fig. 1(a), where the red and blue arrows indicate the spins on the A and B sub-lattices for the ground state. A Möbius strip can then be constructed based on either Fig. 1(b) or Fig. 1(c), which are topologically distinct and satisfy different TBCs to be specified later. Excluding any geometric effects originating from the curvature of the strip [21, 22, 23, 24, 25, 26], we consider a minimal Hamiltonian that preserves the local rotational symmetry about the local $z$ axis:
@@
where $i$ labels the lattice on the strip, $J_{F}$ ($J_{AF}$) is the nearest-neighbor exchange interaction for (between) the same (different) spin species, and $K$ is the perpendicular easy-axis anisotropy. In our convention, all these parameters are positive. Contrary to a single ferromagnet spin chain arranged on a Möbius strip [23, 27], our system is free from geometrical frustration, thus no domain walls are present in the ground state.
To derive the quantum magnon excitations, we apply the linearized Holstein–Primakoff transformations on the spin raising and lowering operators, $S^{\pm}=S_{x}\pm\mathrm{i}S_{y}$, for each sublattice
@@
We cannot directly apply Fourier transformations to Eq. (3) because the PBCs, $a_{i+N}=a_{i}$ and $b_{i+N}=b_{i}$, are explicitly broken. Instead, we have
@@
which means the definitions of A and B chains are interchanged after winding around the Möbius strip by $2\pi$. Regarding Eq. (4), we recombine $a_{i}$ and $b_{i}$ to define the following operators:
@@
where $x_{i}$ specifies the position of site $i$ from $1$ to $N$ along the strip [see Fig. 1], $L$ is the total length of the nano-ribbon, and $\xi=\pm 1$ corresponds the two distinct ways of connection illustrated in Fig. 1(b) and 1(c). The new magnon operators $\alpha_{i}$ and $\beta_{i}$ satisfy not only the bosonic commutation relations but also the PBCs: $\alpha_{i+N}=\alpha_{i}$ and $\beta_{i+N}=\beta_{i}$. Using this new set of basis, the magnon Hamiltonian Eq. (3) becomes
@@
Owing to the bosonic commutation relations of the BdG basis, we need to diagonalize $\sigma_{z}\mathcal{H}_{\alpha(\beta)}$ rather than $\mathcal{H}_{\alpha(\beta)}$ for the magnon solutions [28]. In this regard, we obtain the eigen-frequencies (we set $\hbar=1$)
@@
with the corresponding eigenvectors
$v_{\alpha(\beta)}^{\pm}=$ $\left(\sqrt{q_{\alpha(\beta)}^{1}}\pm\sqrt{q_{\alpha(\beta)}^{2}},\sqrt{q_{\alpha(\beta)}^{1}}\mp\sqrt{q_{\alpha(\beta)}^{2}}\right)^{\mathrm{T}}$ (11)
@@
To better understand the magnon eigenmodes, we now express Eq. (12) in terms of the original spin variables. Using $S^{\pm}=S^{x}\pm\mathrm{i}S^{y}$ and Eqs. (2) and (7), we obtain
@@
$\qquad\qquad\qquad\ \left.+\sqrt{q_{\beta}^{1}}S_{Bi}^{x}+\mathrm{i}\sqrt{q_{\beta}^{2}}S_{Bi}^{y}\right],$ (13b)
where it is important to note that $q_{\alpha(\beta)}^{1}>q_{\alpha(\beta)}^{2}>0$. In the classical limit, both the $\alpha$ and $\beta$ branches feature a right-handed (left-handed) elliptical precession of $\bm{S}_{A}$ ($\bm{S}_{B}$) with the major axes lying in the $x$ direction; $\bm{S}_{A}$ and $\bm{S}_{B}$ always precess about the easy-axis with the same amplitude and opposite chirality. The distinctions between the two branches manifest in two aspects, as schematically demonstrated in Fig. 2(a)–(c). First, from a local bird-eye view, $\bm{S}_{A}$ and $\bm{S}_{B}$ in the $\alpha$-mode at $k=0$ overlap with each other when passing the minor axes of their elliptical trajectories while becoming back-to-back when passing the major axes. On the contrary, $\bm{S}_{A}$ and $\bm{S}_{B}$ in the $\beta$-mode at $k=0$ are back-to-back on the minor axes while overlapping each other on the major axes. Second, due to the momentum shift $\delta k=\pi/L$ in the $\alpha$ branch, the spin precessions on site $i=N$ differ from those on site $i=1$ by a $\pi$ phase even for $k=0$, which exactly compensates the impact of the Möbius TBC that connects sites $1$ and $N$ with a flip. In contrast, the $\beta$ mode at $k=0$ does not exhibit a phase difference between $i=1$ and $i=N$, which, in combination with the first distinction above, is just commensurate with the Möbius TBC.
Discussion. To demonstrate the unique characteristics of the magnon eigenmodes, we draw a 3D perspective in Fig. 2(c) where $\bm{S}_{A}$ and $\bm{S}_{B}$ share the same origin such that their precessional trajectories are concentric about the local $z$ axis. It is easy to deduce that the Néel vector $\bm{n}=(\bm{S}_{A}-\bm{S}_{B})/2S$ undergoes a pendulum-like oscillation restricted to the plane containing the major (minor) axes of the two elliptical trajectories in an $\alpha$ mode ($\beta$ mode). Comparatively, the total spin vector $\bm{S}=(\bm{S}_{A}+\bm{S}_{B})/2S$ oscillates linearly on a plane orthogonal to that of the Néel vector. That is to say, by the standard of spin wave polarization [19, 20], both the $\alpha$ and $\beta$ modes are linearly polarized, thus being devoid of chirality.
The above intuitive picture can be corroborated by a straightforward algebra. According to Eqs. (12) and (13), the real-time evolution of the classical spin vectors is
@@
for the $\alpha$- and $\beta$-branch, respectively, where the $+$ (--) sign corresponds to the A (B) sublattice. From Eq. (14), we can read off the oscillating components of the Néel vector and the total spin vector as

### v4

# Topology-induced symmetry breaking demonstrated in antiferromagnetic magnons on a Möbius strip

@@ Abstract
We propose a mechanism of topology-induced symmetry breaking, where certain local symmetry preserved by the Hamiltonian is explicitly broken in the eigenmodes of excitation due to topologically nontrivial boundary conditions. We demonstrate this phenomenon by studying magnonic excitations on a Möbius strip comprising of two antiferromagnetically coupled spin chains. Even with a simple Hamiltonian respecting local rotational symmetry and all local curvature effects ignored, magnons exhibit linear polarization of the Néel vector devoid of chirality, forming two non-degenerate branches that cannot be smoothly connected to or be decomposed by the circularly-polarized magnons. Correspondingly, one branch undergoes a spectral shift and only admits standing waves of half-integer wavelength, whereas the other only affords standing waves of integer wavelength. The suppression of chirality in the order parameter dynamics, hence the pattern of standing waves, can be generalized to other elementary excitations on non-orientable surfaces. Under the Möbius boundary condition, we further identify an exotic phase hosting spontaneous antiferromagnetic order whilst all exchange couplings are ferromagnetic. Our findings showcase the profound influence of real-space topology on the physical nature of not just the ground state but also the quasiparticles.
I Introduction
I.1 Background
@@
In solid-state systems, symmetry and interactions can directly manifest in the physical properties of quasiparticles (i.e., quanta of elementary excitations) by affecting the momentum-space topology, whereas the subtle impact of real-space topology remains elusive. Prevailing studies customarily adopt periodic boundary conditions (PBCs) in the real space when dealing with quasiparticles [1], for which the system becomes topologically equivalent to a circle, a torus, or a 3D-torus depending on the dimensionality. However, there exist exotic structures such as Möbius strips that inherently do not conform with the PBCs and cannot be smoothly deformed into tori. Figure 1 illustrates that a Möbius strip is non-orientable by nature as it consists of a single surface and a single edge, which leads to an ambiguous global normal vector, precluding the validity of ordinary PBCs.
@@
Concerning the physical behavior of quasiparticles on such a non-orientable object and their subtle relations with the ground state, it is tempting to ask: what are the implications of the topologically non-trivial boundary conditions? Recently, quasiparticles residing on Möbius strips aroused increasing theoretical attentions [2, 3, 4, 5, 6, 7, 8, 9]. On the experimental side, Möbius strips have been realized in a wide variety of systems such as molecules [10, 11], single crystals [12], resonators [13, 14, 15], and optical cavities [16, 17], fertilizing a vibrant arena for exploring new physics emerging from the Möbius topology. Nevertheless, most of these studies focused on the local curvature effects under continuous geometry. It is far from clear if there exist any residual consequences arising only from the Möbius boundary condition when spatial curvature is discarded, which can even survive in the limit of very large systems.
Figure 1: A Möbius strip is non-orientable because the normal vector cannot be globally defined and the PBCs are inapplicable. As a result, chirality associated with the normal vector (illustrated in dashed arrows) becomes ambiguous, rendering the chiral modes of elementary excitations (such as circularly polarized magnons) at odds with the boundary condition.
I.2 General Considerations
A direct consequence following the ambiguity of global normal vector is the absence of a globally defined chirality. As illustrated in Fig. 1, the right-handed direction locally associated with the normal vector inevitably conflicts with itself as we loop around the Möbius strip for a full cycle, fundamentally disrupting the eigenmodes of excitations of all kinds accompanied by chirality. In particular, it is impossible to distinguish between the right-handed and left-handed polarizations.
To be more quantitative, let us construct a local orthogonal basis shown in Fig. 1, wherein $z$ coincides with the local normal vector while $x$ ($y$) is along the longitudinal (transverse) direction of the strip. We represent the linearly polarized modes along $x$ and $y$ by |$\ket{\phi_{x}(\bm{r})}$ and |$\ket{\phi_{y}(\bm{r})}$, respectively, with $\bm{r}=\{x,y\}$ specifying the location on the strip. Through gauge fixing, |$\ket{\phi_{x}(\bm{r})}$ and |$\ket{\phi_{y}(\bm{r})}$ can be locally in phase. Generally, a mode with chirality is expressed as a coherent superposition:
|$\ket{\psi}=\sum_{\bm{r}}\mathcal{N}(\bm{r})\left[\ket{\phi_{x}(\bm{r})}+f(\bm{r})e^{\mathrm{i}\theta(\bm{r})}\ket{\phi_{y}(\bm{r})}\right],$ (1)
where $\mathcal{N}(\bm{r})$ is a normalization factor; $f(\bm{r})$ and $\theta(\bm{r})$ are independent real continuous functions. At $\bm{r}$, |$\ket{\psi}$ exhibits the right-handed (left-handed) chirality for $\theta(\bm{r})<0$ ($>0$). The corresponding classical trajectory is elliptical for $\theta(\bm{r})\neq\pm\pi/2$, with the principal axes along $\tan^{-1}f(\bm{r})$ and $\tan^{-1}f(\bm{r})+\pi/2$. When $\theta(\bm{r})=\pm\pi/2$, the trajectory is circular and all directions can be viewed as principal axes. When $\theta(\bm{r})\rightarrow 0$, the trajectory shrinks to a line so that the polarization becomes linear. In the summation over $\bm{r}$, $x$ runs from 00 to $L$ with $L$ labeling the length of the Möbius strip, while $y$ runs over the width from $-w/2$ to $+w/2$. Should $\theta(\bm{r})$ be allowed to keep the same sign everywhere on the strip, the mode represented by |$\ket{\psi}$ will acquire a globally well-defined chirality.
To impose the boundary conditions, we define $\hat{T}_{L}$ as the translation operator (along $x$) that moves functions of $\bm{r}$ from $(x,y)$ to $(x+L,y)$ by continuously sliding the local coordinate frame along the strip. Since the linearly polarized modes are not entitled with chirality in priori and are not disrupted by the ambiguity of the normal vector, they must be single valued, thus
$\hat{T}_{L}\ket{\phi_{x}(x,y)}=\ket{\phi_{x}(x+L,y)}=\ket{\phi_{x}(x,-y)},$ (2a)
$\hat{T}_{L}\ket{\phi_{y}(x,y)}=\ket{\phi_{y}(x+L,y)}=\ket{\phi_{y}(x,-y)}.$ (2b)
On the other hand, the superposition coefficients satisfy
$\hat{T}_{L}\mathcal{N}(x,y)=\mathcal{N}(x+L,y)=\mathcal{N}(x,-y),$ (3a)
$\hat{T}_{L}f(x,y)=f(x+L,y)=f(x,-y),$ (3b)
$\hat{T}_{L}\theta(x,y)=\theta(x+L,y)=-\theta(x,-y),$ (3c)
where $-1$ in the last equation is interpreted as $e^{\pm\mathrm{i}\pi}$ with $\pm$ referring to the two distinct ways of wrapping up the Möbius strip [see the schematics in Fig. 2(b) and (c)]. In general, if a strip is multiply twisted when attaching its two ends, $\hat{T}_{L}\theta[x,y]=\exp(2\mathrm{i}p\pi)\theta[x,(-1)^{2|p|}y]$ where $p$ registers the number of turns on the strip. The Möbius strip is half-twisted so $p=\pm 1/2$.
If |$\ket{\psi}$ serves as an eigenmode of excitation, it must be single valued, i.e., invariant after traversing the Möbius strip by $L$, or simply $\hat{T}_{L}\ket{\psi}=\ket{\psi}$. Applying the above boundary conditions, we obtain
$\hat{T}_{L}\ket{\psi}$ $=\sum_{\bm{r}}\mathcal{N}(x,-y)\times$
[|$\left[\ket{\phi_{x}(x,-y)}+f(x,-y)e^{-\mathrm{i}\theta(x,-y)}\ket{\phi_{y}(x,-y)}\right]$
$=$ $\sum_{\bm{r}^{\prime}}\mathcal{N}(\bm{r}^{\prime})\left[\ket{\phi_{x}(\bm{r}^{\prime})}+f(\bm{r}^{\prime})e^{-\mathrm{i}\theta(\bm{r}^{\prime})}\ket{\phi_{y}(\bm{r}^{\prime})}\right],$ (4)
where the symmetric range of summation over $y$ has been exercised when $\bm{r}^{\prime}$ replaces $\bm{r}$ in the last step. Comparing Eq. (4) with Eq. (1), we have $\hat{T}_{L}\ket{\psi}=\ket{\psi}$ if and only if $\theta(\bm{r})$ vanishes identically, indicating that |$\ket{\psi}$ is strictly linearly-polarized. That is to say, a single-valued mode on the Möbius strip cannot possess a definite chirality, whereas a mode with prescribed chirality cannot be single valued, let alone being an eigenmode.
The Möbius topology does not place a strong restriction on the Hamiltonian $H$, so long as $H$ is single valued and respects $\hat{T}_{L}H(x,y)=H(x+L,y)=H(x,-y)$. For instance, the Heisenberg Hamiltonian acting only in the spin space preserves the local spin-rotational symmetry, not contradicting with the boundary conditions. In the topological trivial scenarios, this symmetry ensures the existence of circularly-polarized eigenmodes in the spin excitations, which are furnished by well-defined chirality. On the other hand, the Möbius topology can drastically modify these eigenmodes and suppress their chirality, as analyzed above. At this point, it is inspiring to conceive a hitherto overlooked scenario:
Certain local symmetry preserved by the Hamiltonian is explicitly broken in the eigenmodes of excitation (or quasiparticles) owing to the real-space topology.
We name the underlying mechanism topology-induced symmetry breaking (TISB), which unravels the extraordinary behavior of quasiparticles exclusively enabled by the topological boundary conditions (TBCs) but has nothing to do with the local curvature effect. While TISB could entail profound consequences in various physical systems with non-trivial real-space topology, in this paper we demonstrate its manifestation in the magnonic excitations on a Möbius strip.
I.3 System Description
To avoid confusion with the well-known mechanism of spontaneous symmetry breaking associated with the ground state, here we consider a special case where the ground state remains intact and trivial, whereas the eigenmodes of excitation are fundamentally disrupted by the Möbius topology. As illustrated in Fig. 2(a), we consider a nano-ribbon composed of two ferromagnetic spin chains that are oppositely aligned, forming an effective antiferromagnetic (AFM) system. The nano-ribbon can form a Möbius strip in two distinct ways according to Fig. 2(b) and (c). To motivate our study, we first make a critical observation: although the AFM configuration in the ground state is compatible with the Möbius TBC, the magnon excitations with circular polarizations are inherently irreconcilable with the system topology. To be specific, when only the exchange interactions and the easy-axis anisotropy are considered, the spin Hamiltonian respects the local $O(2)$ rotational symmetry in the spin space. Consequently, if the TBC is disregarded, the magnon eigenmodes would become circularly polarized, exhibiting either left-handed or right-handed chirality for both spin species [18, 19, 20, 21, 22]. For example, in the left-handed mode depicted in Fig. 2(a), $\bm{S}_{A}$ and $\bm{S}_{B}$ both rotate clockwisely but $\bm{S}_{A}$ has a larger oscillation amplitude. However, as illustrated in Fig. 2(b) and (c), imposing the Möbius TBC by connecting the $1$ and $N$ sites (with A and B swapped) will inevitably disrupt both the chirality and the amplitude of the spin precessions. Therefore, the magnon solutions that are commensurate with the Möbius topology must be fundamentally different from the circularly-polarized modes widely known for collinear AFM materials.
@@ I.3 System Description
II Model and Solutions
Consider a nano-ribbon with the AFM configuration shown in Fig. 2(a), where the red and blue arrows indicate the spins on the A and B sub-lattices in the ground state. A Möbius strip can then be constructed based on either Fig. 2(b) or Fig. 2(c), which are topologically distinct and satisfy different TBCs to be specified later. Excluding the geometric effects originating from the curvature of the strip [23, 24, 25, 26, 27, 28, 29], we adopt a minimal Hamiltonian that preserves the local rotational symmetry about the local $z$ axis:
@@ II Model and Solutions
where $i$ labels the lattice on the strip, $J_{F}$ ($J_{AF}$) is the nearest-neighbor exchange interaction for (between) the same (different) spin species, and $K$ is the perpendicular easy-axis anisotropy. The collinear AFM ordering depicted in Fig. 2 is well protected by the magnetic anisotropy $K$, which suppresses quantum fluctuations in the reduced dimension. The summation $\langle i,j\rangle$ includes the link between 1 and $N$ such that everywhere on the strip looks the same regardless of the numbering. In our convention, all parameters are positive. Contrary to a single ferromagnetic spin chain arranged on a Möbius strip [23, 24], our system is free from geometrical frustration, thus no domain wall is present.
To derive the quantum magnon excitations, we apply the linearized Holstein–Primakoff transformations on the spin raising and lowering operators (assuming that $S$ is sufficiently large), $S^{\pm}=S^{x}\pm\mathrm{i}S^{y}$, for each sublattice
@@
We cannot directly apply Fourier transformations to Eq. (7) because the PBCs, $a_{i+N}=a_{i}$ and $b_{i+N}=b_{i}$, are explicitly broken. Instead, we have the TBCs
@@
which means the definitions of A and B chains are interchanged after winding around the Möbius strip. Basing on Eq. (8), we can recombine $a_{i}$ and $b_{i}$ as
@@
where $x_{i}$ specifies the position of site $i$ from $1$ to $N$ along the strip [see Fig. 2], $L$ is the total length of the nano-ribbon, and $\xi=\pm 1$ corresponds to the two distinct ways of connection illustrated in Fig. 2(b) and 2(c). The new magnon operators $\alpha_{i}$ and $\beta_{i}$ satisfy not only the bosonic commutation relations but also the PBCs
$\alpha_{i+N}=\alpha_{i},\qquad\beta_{i+N}=\beta_{i}.$ (10)
Under this new set of basis, Equation (7) becomes
@@
Owing to the bosonic commutation relations of the BdG basis, we need to diagonalize $\sigma_{z}\mathcal{H}_{\alpha(\beta)}$ rather than $\mathcal{H}_{\alpha(\beta)}$ for the magnon solutions [30, 31]. In this regard, we obtain the eigen-frequencies (we set $\hbar=1$)
@@
with the corresponding eigenvectors (unnormalized)
$v_{\alpha(\beta)}^{\pm}=$ $\left(\sqrt{q_{\alpha(\beta)}^{1}}\pm\sqrt{q_{\alpha(\beta)}^{2}},\sqrt{q_{\alpha(\beta)}^{1}}\mp\sqrt{q_{\alpha(\beta)}^{2}}\right)^{\mathrm{T}},$ (16)
@@
To intuitively understand the magnon eigenmodes, we now express Eq. (17) in terms of the original spin variables. Using Eqs. (6) and (12), we obtain
@@
$\qquad\qquad\qquad\ \left.+\sqrt{q_{\beta}^{1}}S_{Bi}^{x}+\mathrm{i}\sqrt{q_{\beta}^{2}}S_{Bi}^{y}\right].$ (18b)
The classical limit of Eqs. (6) can be established by considering the quantum averages: $\langle S_{A}^{-}\rangle=\sqrt{2S}\langle a^{\dagger}\rangle$ and $\langle S_{B}^{+}\rangle=\sqrt{2S}\langle b^{\dagger}\rangle$ correspond to the left-handed precessions of $\bm{S}_{A}$ and $\bm{S}_{B}$ with respect to the equilibrium direction of $\bm{S}_{A}$ (the reference direction). By a straightforward algebra, we obtain the real-time evolution of the classical spin vectors as
@@
for the $\alpha$- and $\beta$-branch, respectively, where the $+$ (--) sign corresponds to the A (B) sublattice. Because $q_{\alpha(\beta)}^{1}>q_{\alpha(\beta)}^{2}>0$, both the $\alpha$ and $\beta$ branches exhibit left-handed (right-handed) elliptical precession of $\bm{S}_{A}$ ($\bm{S}_{B}$) with the major axes of the trajectory lying along $x$; $\bm{S}_{A}$ and $\bm{S}_{B}$ always precess about the easy-axis with the same amplitude but opposite chirality. As a result, the Néel vector $\bm{n}=(\bm{S}_{A}-\bm{S}_{B})/2S$ exhibits linear oscillation devoid of chirality, on which we will further elaborate in the next section.
The distinctions between the two branches manifest in two aspects, as demonstrated in Fig. 3(a)–(b). First, from a local bird-eye view, $\bm{S}_{A}$ and $\bm{S}_{B}$ in the $\alpha$-mode at $k=0$ overlap with each other when passing the minor axes of their elliptical trajectories while becoming back-to-back when passing the major axes. On the contrary, $\bm{S}_{A}$ and $\bm{S}_{B}$ in the $\beta$-mode at $k=0$ are back-to-back on the minor axes while overlapping on the major axes. Second, due to the momentum shift $\delta k=\pi/L$ in the $\alpha$ branch, the spin precessions on site $i=N$ differ from those on $i=1$ by a $\pi$ phase even for $k=0$, which exactly compensates the impact of the Möbius TBC that connects sites $1$ and $N$ with a half twist. In contrast, the $\beta$ mode at $k=0$ does not exhibit a phase difference between $1$ and $N$, which, in combination with the first distinction, is just commensurate with the TBC.
III Analysis and Discussion
To further understand the unique characteristics of the magnon eigenmodes obtained above, we draw a 3D perspective in Fig. 3(c) where $\bm{S}_{A}$ and $\bm{S}_{B}$ share the same origin such that their precessional trajectories are concentric about the local $z$ axis. It is easy to deduce that the Néel vector $\bm{n}=(\bm{S}_{A}-\bm{S}_{B})/2S$ undergoes a pendulum-like oscillation confined on the plane containing the major (minor) axes of the two elliptical trajectories in an $\alpha$ mode ($\beta$ mode). Comparatively, the total spin vector $\bm{S}=(\bm{S}_{A}+\bm{S}_{B})/2S$ oscillates linearly on a plane orthogonal to that of $\bm{n}$. Per the definition of spin wave polarization (in terms of the order parameter dynamics) [19, 20, 21], both the $\alpha$ and $\beta$ modes are linearly polarized, thus being devoid of chirality.
The above intuitive picture can be corroborated by Eq. (19), from which we can further read off the oscillating components of the Néel vector and the total spin vector as